# Consumption Partial Insurance

### v3

@@ 5.3 Quadratic consumption function
We now turn to the estimation of the quadratic consumption function. The results appear in column 5 of table 4; the full set of targeted moments is shown in appendix B.4.3.353535Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 3., [^36] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case, i.e., in column 4, table 4. This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model. This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2). However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway (section 3). Our estimates of the linear function in that case do not differ from those in column 4, table 4, so estimation is robust to the choice of moments. These results are available upon request.
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, highly significant and close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial in the shocks. We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks (statistically significant at the margin for now), while we estimate a positive $\psi^{(2)}=0.015$ for transitory shocks (statistically insignificant for now). The coefficient $\omega^{(22)}$ on the product of permanent and transitory shocks is large but imprecisely estimated. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^37] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
@@
First, bad (negative) permanent shocks are more hurtful than good (positive) ones because they are associated with a much larger transmission parameter into consumption. Consider, for example, a large permanent shock to household income, $\zeta_{it}=\pm 0.5$, approximately three standard deviations away from the mean. If the shock is negative ($\zeta_{it}=-0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$; by contrast, if the shock is positive ($\zeta_{it}=0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$. We can generalize this to bad and good shocks of any magnitude and conclude that
@@
Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 3, but they also have much larger pass-through rates into consumption. Bad permanent shocks are thus double hurtful.
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$. We can generalize this to bad shocks of any magnitude and conclude that
@@
This result is not hard to comprehend. When a bad shock hits, households can insure consumption by running down savings or by using other formal or informal insurance arrangements. But there is clearly a limit to this. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, thus resulting in a larger fraction of the shock eventually transmitting into consumption. [^38] The parameter estimates also imply that the pass-through of a good shock decreases with its magnitude. While this seems counterintuitive, recall that extreme good shocks are rare in our data, which follows from left skewness and excess kurtosis in table 3. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios. For example, Bunn et al. (2018) find that the marginal propensity to consume from negative income shocks is larger than from positive shocks, while Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the shock. This literature, however, elicits responses to transitory shocks. By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
Why would a positive or small shock be better insured? While we cannot assess candidate mechanisms without a structural model, our best explanation is that people save a good income shock (thus consumption does not respond much) but they dissave in the presence of a bad one. As there is clearly a limit to how much households can dissave or borrow, especially in the presence of borrowing constraints, a large negative shock would typically affect consumption by relatively more. [^39] Busch and Ludwig (2021) simulate a lifecycle model with concave utility and borrowing constraints; they find larger pass-through of negative versus positive permanent shocks. While we do not model borrowing constraints here, a strength of our approach is that it enables us to measure partial insurance in the data without taking a stance on the precise underlying structural model/economic mechanism at play.
@@
$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.$
@@
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption. This is similar to our findings for the linear case as well as to BPP. By contrast, this is not the case for certain well-defined subsamples, but we postpone the discussion of the subsamples for section 5.4.
@@
Despite the lack of statistical significance, the partial derivative with respect to transitory shocks has, at face value, an interesting implication: large good transitory shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero. While we should be cautious in drawing conclusions about extreme good transitory shocks (given left skewness and excess kurtosis, large good transitory shocks are not present in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and other semi-structural studies that use covariance restrictions on survey data. The former studies typically exploit large and infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021). In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions. Our nonlinear consumption function reconciles the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
@@
To summarize our findings from the quadratic consumption function: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do. For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein. By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way. [^40] Appendix table C.7 shows that similar conclusions are obtained if we use consumption imputed from the CEX. In particular, $\phi^{(2)}<0$ (albeit not statistically significant), with a magnitude similar to that obtained directly from the PSID ($-0.036$ versus $-0.04$), and $\psi^{(2)}>0$.
5.4 Sensitivity to age, wealth, and education
@@
We first consider how partial insurance changes with age, by repeating the estimation over three, almost equally sized, subsamples of households with the male spouse between 30–40, 41–50, and 51–65 years old respectively. [^41] We target the same moments as earlier, broken down by the age of the male spouse. The moments of income shocks are also age-specific but we do not report those results for brevity. Intuitively, the transmission parameters of shocks should exhibit some age dependence since wealth accumulates over the lifecycle, which in turn influences the household’s ability to insure against income fluctuations.
The results appear in table 5. Column 1 presents results from the linear specification when only second moments are targeted. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are included in the estimation. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^42] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
@@
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID
(1) (2) (3) (4)
$\phi^{(1)}$ 30-40 0.154 (0.045) 0.164 (0.051) 0.089 (0.058) 0.149 (0.050)
41-50 0.126 (0.051) 0.146 (0.055) 0.079 (0.089) 0.115 (0.049)
51-65 0.180 (0.049) 0.175 (0.051) 0.136 (0.061) 0.113 (0.057)
$\psi^{(1)}$ 30-40 0.094 (0.079) 0.083 (0.073) 0.095 (0.107) 0.105 (0.090)
41-50 0.010 (0.079) -0.016 (0.072) -0.075 (0.061) 0.019 (0.090)
51-65 -0.046 (0.070) -0.044 (0.066) 0.037 (0.101) -0.063 (0.088)
$\phi^{(2)}$ 30-40 0.006 (0.040)
41-50 -0.008 (0.063)
51-65 -0.105 (0.054)
$\psi^{(2)}$ 30-40 0.036 (0.061)
41-50 0.048 (0.046)
51-65 -0.052 (0.079)
$\omega^{(22)}$ 30-40 -1.090 (1.465)
41-50 -0.290 (1.195)
51-65 3.003 (1.306)
$\sigma^{2}_{\xi}$ 30-40 0.014 (0.002) 0.014 (0.002) 0.016 (0.002) 0.013 (0.004)
41-50 0.022 (0.002) 0.022 (0.002) 0.022 (0.002) 0.022 (0.002)
51-65 0.021 (0.002) 0.021 (0.002) 0.023 (0.002) 0.010 (0.008)
$\gamma_{\xi}$ 30-40 -0.001 (0.001) -0.001 (0.001)
41-50 0.006 (0.001) 0.006 (0.002)
51-65 -0.006 (0.002) -0.006 (0.002)
$\kappa_{\xi}$ 30-40 0.004 (0.002)
41-50 0.007 (0.002)
51-65 0.006 (0.002)
$\sigma^{2}_{u_{c}}$ 30-40 0.046 (0.003) 0.046 (0.003) 0.046 (0.003) 0.046 (0.003)
41-50 0.040 (0.002) 0.040 (0.002) 0.040 (0.002) 0.040 (0.002)
51-65 0.043 (0.002) 0.043 (0.002) 0.043 (0.002) 0.043 (0.002)
$\gamma_{u_{c}}$ 30-40 0.000 (0.002) 0.000 (0.002)
41-50 0.002 (0.002) 0.002 (0.002)
51-65 0.004 (0.002) 0.004 (0.002)
$\kappa_{u_{c}}$ 30-40 0.014 (0.003)
41-50 0.012 (0.002)
51-65 0.012 (0.002)
@@
Notes: The table presents the estimates of the parameters of the consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse (the underlying income moments also vary over these age brackets). Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
Column 4 presents the results from the quadratic specification. In this case, $\phi^{(1)}$ decreases monotonically with age, so the average permanent shock (the shock close to the mean of the distribution) impacts consumption gradually less as households age. In line with theory, this pattern reflects the age-increasing role of financial wealth for self-insurance. Perhaps more importantly, $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases monotonically with age. $\phi^{(2)}$ drops from $0.006$ to $-0.008$ as we move from ages 30–40 to ages 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.105$ in the 51–65 age bracket. As a result, large negative permanent shocks impact consumption more severely at older than at younger ages.
The last finding echoes Guvenen et al. (2021), who show that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the low-earners or the young. In fact, we also find larger left skewness in income growth as households age in the PSID and more negative skewness in permanent shocks among the old. [^43] The coefficient of skewness of permanent shocks more than triples in magnitude from $-0.65$ to $-2.39$ as we move from ages 30–40 to 51–65. The coefficient of skewness of transitory shocks, by contrast, does not change much with the age – its estimate is $-1.43$ and $-1.45$, respectively, over ages 30–40 and 51–65. So, not only do permanent shocks become more negative as households age, but they also transmit into consumption at increasingly larger rates. In other words, not only is there more room for income to fall among older earners, but this also causes even greater harm to consumption. By contrast, despite a similar pattern for transitory shocks, their transmission parameters remain statistically indistinguishable from zero over the entire lifecycle.

### v4

@@ 5.3 Quadratic consumption function
We now turn to the estimation of the quadratic consumption function. The results appear in column 5 of table 3; the full set of targeted moments is shown in appendix B.4.3.353535Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 2., [^36] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case, i.e., in column 4, table 3. This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model. This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2). However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway. Our estimates of the linear function in that case do not differ from those in column 4, table 3, so estimation is robust to the choice of moments. These results are available upon request.
5.3.1 Main results
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial. We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks and a positive (but statistically insignificant for now) $\psi^{(2)}=0.015$ for transitory shocks. The coefficient $\omega^{(22)}$ on the product of the two shocks is large but very imprecisely estimated. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^37] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
@@ 5.3.1 Main results
First, bad (negative) permanent shocks are more unsettling than good (positive) ones because they are associated with a larger transmission parameter into consumption. Consider, for example, a large permanent shock to household income, $|\zeta_{it}|=0.5$, about three standard deviations away from the mean. If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.$=0.036$, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.$=0.036$). We can generalize this to bad and good shocks of any magnitude and conclude that
@@
Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption. Bad permanent shocks are thus double hurtful.
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$). We can generalize this to bad shocks of any magnitude and conclude that
@@
Why would a positive or small shock be better insured? When a good shock hits, households save part of it and consumption does not respond too much. When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements. But there is clearly a limit to how much they can dissave or borrow. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption. A large negative shock thus affects consumption by relatively more than a small or positive shock does. [^38] The results also imply that the pass-through of a good shock decreases with its magnitude. The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks. However, large left skewness and excess kurtosis imply that extreme good shocks are rare. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios. For example, Bunn et al. (2018) find that the marginal propensity to consume from negative income shocks is larger than from positive shocks, while Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the bad shock. This literature, however, elicits responses to transitory shocks. By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
@@
$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\$
@@
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption. This is similar to our findings for the linear case as well as to BPP. By contrast, this is not the case for certain well-defined subsamples, which we analyze in the next subsections.
@@
Despite the lack of statistical significance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero. While we should be cautious in drawing conclusions about large good transitory shocks (given left skewness and excess kurtosis, such shocks are rare in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and similar studies that employ covariance restrictions on survey data. The former typically exploit large infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021). In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions. Our nonlinear consumption function reconciles the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
@@
To summarize our findings from the quadratic consumption function: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do. For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein. By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way. [^39] Similar conclusions are obtained if we use imputed consumption from the CEX. In particular, $\phi^{(2)}<0$, with a magnitude similar to that obtained from the PSID, and $\psi^{(2)}\approx>0$ (appendix table C.8).
5.3.2 Results by age and position in the income distribution
@@
We now consider how partial insurance changes with age, by repeating the estimation over three, almost equally sized, subsamples of households with the male spouse between 30–40, 41–50, and 51–65 years old respectively. [^40] We target the same moments as earlier, broken down by the age of the male spouse. The moments of income shocks are also age-specific (appendix table C.9). Intuitively, the transmission parameters of shocks should exhibit some age dependence since wealth accumulates over the lifecycle, which in turn influences the household’s ability to insure against income fluctuations.
Table 4 reports results from the quadratic specification (see table C.10 for the linear specification). $\phi^{(1)}$ decreases with age, so the average permanent shock impacts consumption less as households age. In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle. $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age. It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket. As a result, large negative permanent shocks impact consumption more severely at older ages. Consider, again, $\zeta_{it}=-0.5$. At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
@@
Consumption fn.: Quadratic
Age household head: 30–40 41–50 51–65
Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID PSID
(1) (2) (3)
$\phi^{(1)}$ 0.149 (0.050) 0.115 (0.049) 0.113 (0.057)
$\psi^{(1)}$ 0.105 (0.090) 0.019 (0.090) -0.063 (0.085)
$\phi^{(2)}$ 0.006 (0.038) -0.008 (0.063) -0.104 (0.052)
$\psi^{(2)}$ 0.036 (0.061) 0.048 (0.046) -0.052 (0.073)
$\omega^{(22)}$1.089 (1.465) -0.294 (1.195) 2.984 (0.791)
$\sigma^{2}_{\xi}$ 0.013 (0.004) 0.022 (0.002) 0.010 (0.005)
$\sigma^{2}_{u_{c}}$ 0.046 (0.003) 0.040 (0.002) 0.043 (0.002)
@@
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the household head (male spouse). The underlying income moments also vary over these age brackets (appendix table C.9). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. Linear specification estimates appear in appendix table C.10.
Guvenen et al. (2021) document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the young. In our data, skewness of permanent shocks becomes more negative as households age; it more than triples in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9). The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption technology itself changes ($\phi^{(2)}$ more negative). The type of events that underlie large income downturns at older ages – such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption technology that differs between the old and the young. This does not mean that moderate bad shocks are worse insured among the old; for small shocks, the decreasing pattern in $\phi^{(1)}$ more than counteracts the increasing magnitude of $|\phi^{(2)}|$.
Finally, the transmission parameters of transitory shocks remain statistically indistinguishable from zero over the entire lifecycle.
We next consider how partial insurance changes with the household’s position in the income distribution before the shocks hit at $t$, i.e., in the distribution of $y_{it-1}$. Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, unlike ABB whose nonlinear income process naturally allows this. We thus take an empirical approach and estimate the model over different cuts of the distribution of $y_{it-1}$. Sample size considerations prevent us from being overly granular; we tried estimation over deciles, quartiles, and tertiles, and the latter was the most robust. For brevity we report results in the bottom (‘poor’) and top (‘rich’) tertiles. [^41] We target the same moments as earlier, broken down by tertiles of the distribution of past income $y_{it-1}$. The moments of shocks are also income level-specific (appendix table C.9).
Table 5 reports results from the quadratic specification (see table C.11 for the linear specification). Three points emerge. First, $\phi^{(1)}$ among the rich is half the value of the poor, so the average permanent shock affects consumption less strongly among the rich. This reflects again the self-insurance role of wealth given that income-richer households are also wealthier. Second, $\phi^{(2)}$ is negative, confirming the baseline finding that negative permanent shocks have larger pass-through than equal-sized positive shocks. Third, the transmission of transitory shocks shows mixed patterns but it is never statistically significant.
At face value, $\phi^{(2)}$ is more negative among the poor, suggesting that a bad shock of a given size transmits more strongly among them. While factually correct given the estimates, this provides an incomplete view of the underlying dynamics because the distribution of income risk is vastly different between poor and rich. Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but very negative among the rich, with a coefficient of $-4.86$ (multiple times larger than baseline skewness, and highly significant). One may thus study how large bad/good shocks from a given position in the underlying distribution, for example shocks in the 3rd and 97th percentiles, affect consumption in each case. Simulations are needed to infer the specific shocks. [^42] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9. We provide details on similar simulations in section 6. Adopting this approach, the bad shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) but it shifts $c_{it}$ by $-0.035$, twice as much, among the rich (pass-through $0.084$). By contrast, the good shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but shifts $c_{it}$ by only $0.006$ among the rich (pass-through $0.057$). These findings agree with ABB, who simulate the transmission of 10- and 90-percentile shocks and also find a stronger response to bad news among households in the top end of the income distribution and a stronger response to good news among households in the bottom end. The reported $\Delta c_{it}$ are also numerically in the same ballpark (impulse responses with heterogeneity in their appendix). Our results thus suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
Table 5: Estimates of the consumption function, over position in income distribution
Consumption fn.: Quadratic
Lag income $y_{it-1}$: Bottom tertile Top tertile
Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID
(1) (2)
$\phi^{(1)}$ 0.113 (0.029) 0.063 (0.100)
$\psi^{(1)}$ 0.186 (0.199) 0.275 (0.200)
$\phi^{(2)}$0.064 (0.029) -0.025 (0.101)
$\psi^{(2)}$ 0.210 (0.220) -0.268 (0.197)
$\omega^{(22)}$ 1.965 (1.513) 1.072 (2.144)
$\sigma^{2}_{\xi}$ 0.011 (0.005) 0.025 (0.004)
$\sigma^{2}_{u_{c}}$ 0.044 (0.002) 0.041 (0.002)
Notes: The table presents the estimates of the parameters of the quadratic consumption function in the bottom and top tertiles of the distribution of lag income $y_{it-1}$. The underlying income moments also vary over the income distribution (appendix table C.9). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. Linear specification estimates appear in appendix table C.11.
Our approach highlights two distinct forces at play in the response of consumption to income shocks: the underlying nature of income risk (i.e., the income parameters) and the consumption technology (i.e., the consumption parameters). The larger response of the rich to shocks in the bottom of the distribution is not because their consumption technology differs vastly from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates the response to said shocks), but mostly due to stronger downwards risk that the rich face. To infer that, one needs to evaluate the marginal distribution of income risk. ABB’s quantile method can flexibly estimate the consumption response over income quantiles and compare responses to, say, 10- and 90-percentile shocks, but it does not measure the higher-order features of the distribution these shocks come from. This may lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be very similar to one another since the underlying distribution is skewed and leptokurtic. For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked further out at the 3rd and 97th percentiles. Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
$c_{it}=g_{t}(\underbrace{y_{it-1}+F_{t}^{-1}(u_{it})}_{y_{it}},~{}\dots),\$
where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ is a uniform shock denoting the quantile of the shocks’ distribution $F_{t}$. They then simulate the consumption pass-through over different values for $y_{it-1}$ and $u_{it}$, without measuring $F_{t}$. Our model allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the sign and size of the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$. It can thus describe more transparently how income shocks of different signs/sizes shift consumption.
How does a less negative $\phi^{(2)}$ among the rich reconcile with a more negative $\phi^{(2)}$ among the old, who are presumably also rich? The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between the nature of income risk and the consumption technology. Skewness of permanent shocks is twice as large among the rich ($-4.86$) as it is among the old ($-2.39$). The consumption distribution is not too different between the two, however, so the rich must have more insurance to a bad shock of given size. This explains why $\phi^{(2)}$ is four times smaller among them. In both cases (old, rich), bad permanent shocks from a given position in the distribution decrease consumption more strongly than among their counterparts (young, poor).
5.3.3 Results by wealth and education